Simplify the expression 7/24-3/14÷1/3


1
Expert's answer
2021-05-07T09:24:19-0400
"\dfrac{7}{24}-\dfrac{3}{14}\div\dfrac{1}{3}=\dfrac{7}{24}-\dfrac{3}{14}\cdot3"

"=\dfrac{7}{24}-\dfrac{9}{14}=\dfrac{49-108}{168}=-\dfrac{59}{168}"